Baichung's father is 26 years younger than baichung's grandfather and 29 years older than baichung.the sum of all the three is 135 years.what is the age of each one of them?give the answer with calculation
1
Expert's answer
2013-09-12T13:32:47-0400
Solution
If age of Baichung's father is x, age of baichung's grandfather is y=x+26, age of baichung z=x-29, we havex+y+z=3x-3=135 years. From hence
x=46 years,
y=72 years,
z=17 years.